OpenStudy (anonymous):

Simplify 7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54. 11 square root of 6 end root minus 7 square root of 12 11 square root of 3 end root minus 7 square root of 6 negative 3 square root of 9 4 square root of 9

OpenStudy (anonymous):

can someone please help me

OpenStudy (anonymous):

I really don't understand this I need HELP PLEASE!!!!!!

OpenStudy (anonymous):

\[7 \sqrt{3}-4 \sqrt{6}+\sqrt{48}-\sqrt{54} \]\[7 \sqrt{3}-4 \sqrt{6}+4 \sqrt{3}-3 \sqrt{6} \]\[7 \sqrt{3}+4 \sqrt{3}-3 \sqrt{6}-4 \sqrt{6} \]\[11 \sqrt{3}-7 \sqrt{6} \]

OpenStudy (anonymous):

I still don't get this I am new to this

OpenStudy (anonymous):

I will give you a medal I just need help to understand.

OpenStudy (anonymous):

#jaynator495,#Michele_Laino HELP ME